UT March 16 2019

OSA: D.G. Huerta

ISp: J. Morris

ISA: B. Rothberg (Tucson)

AO ARGOS: S. Walsh (Tucson)
ARGOS Commander: G. Rahmer (Tucson)

Observers: B. Rothberg, S. Walsh, G. Rahmer (first part), A. Conrad, O. Kuhn (2nd part).

Primary Goal: ARGOS testing

Backup: MODS-Bino pointing tests


UT 01:29 - LUCI software started up - closed for clouds

UT 02:09 - MODS1&2 awake

UT 02:13 - sievesnap done - all OK

UT 03:03 - still closed - ARGOS giving up - if we open - will do pointing

UT 03:24 - configuring for Bino-MODS

UT 05:05 - still closed

UT 05:30 - opening enclosure - will give it a try

-will do rotator center measurement first, then proceed to pointing stars

UT 05:55 - well.... we were about to start collimating when the clouds further enveloped us....+60 dec is clear, below that....(make that +70)

-closing back up

UT 09:35 - still closed, relief crew has arrived, turning it over to Olga and Al for the TOO....good night and good luck...

-- BarryRothberg - 16 Mar 2019
Topic revision: r3 - 16 Mar 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback